OpenStudy (anonymous):

Simplify: (x-4)(x+3)

OpenStudy (anonymous):

x^2-x-12

OpenStudy (anonymous):

x*x-4*x+x*3-4*3 = x^2-4x+3x-12 = x^2-x-12

OpenStudy (anonymous):

Thank you all